OpenStudy (cydney_morgan):

Is this the correct answer? (Omg I am so sorry to bother you again) Energy levels are known as the regions within the electron cloud where electrons are located. the different types of energy within elements. the distance an electron is from the nucleus. the path an electron follows when orbiting around the nucleus. Is this the correct answer? the regions within the electron cloud where electrons are located. (I am so bad with atoms please forgive me if I am wrong)

OpenStudy (cydney_morgan):

@Abhisar Sorry to tag and spam you! >.<

OpenStudy (cydney_morgan):

Because electrons do not follow one straight path right? Or am I completely wrong.

OpenStudy (abhisar):

y do u say so many sorries :O

OpenStudy (cydney_morgan):

Because I hate to bother people! It makes me feel bad!

OpenStudy (abhisar):

you are not bothering anyone...m here to help only :) \(\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \Large\cdot \quad \cdot\\ \cdot\\ \huge \smile \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\)

OpenStudy (cydney_morgan):

Thank you! :)

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Cydney_Morgan Is this the correct answer? (Omg I am so sorry to bother you again) Energy levels are known as the regions within the electron cloud where electrons are located. \(\huge\checkmark\) the different types of energy within elements. the distance an electron is from the nucleus. the path an electron follows when orbiting around the nucleus. Is this the correct answer? the regions within the electron cloud where electrons are located. (I am so bad with atoms please forgive me if I am wrong) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (cydney_morgan):

Yay! I'm correct!! :D

OpenStudy (abhisar):

Heisenberg's uncertainity principle stated that position of electron can't be calculated accurately, that's y say region wr the electron cloud can be found.

OpenStudy (cydney_morgan):

They travel in waves correct?

OpenStudy (abhisar):

yes ! That's De-broglie's dual nature of matter !

OpenStudy (cydney_morgan):

Yay! :D

OpenStudy (abhisar):

Just like light has both wave and matter nature, electrons also has both wave and matter nature.

OpenStudy (cydney_morgan):

Thank you! (I'm writing this stuff in my notes)

OpenStudy (abhisar):

\(\color{red}{\huge\bigstar}\huge\text{You're Most Welcome! }\color{red}\bigstar\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)